Suppose you have a project that has a 0.8 chance of tripling your investment in a year and a 0.2 chance of halving your investment in a year. What is the standard deviation of the rate of return on this investment? (Do not round intermediate calculations. Enter your answer as a decimal rounded to 4 places.)Suppose you have a project that has a 0.8 chance of tripling your investment in a year and a 0.2 chance of halving your inves
Answer #1

Please refer to below spreadsheet for calculation and answer. Cell reference also provided.

В C D Initial investment (assumed) $100.00 Return Probability 0.8 0.2 Investment value at end of year $300.00 $50.00 200% -50

Cell reference -

2 А в Initial investment (assumed) 100 Probability Investment value at end of year Return 0.8 0.2 =C2*3 =C2/2 =C5/C2-1 =C6/C2

Hope this will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.